Let f(a) = 6 – V22 + 1 and g(2) = 2 – 2. Then (fog)(5) = (fºg)(20) =
Answer #1

\small f(x) = 6-\sqrt{x^{2}+1}\textup{ and }g(x)=x-2

\small f(g(x))=6-\sqrt{(x-2)^{2}+1}=6-\sqrt{x^{2}-4x+4+1}=6-\sqrt{x^{2}-4x+5}

\small \boldsymbol{f(g(x))=6-\sqrt{x^{2}-4x+5}}

\small f(g(5))=6-\sqrt{5^{2}-4(5)+5}

\small f(g(5))=6-\sqrt{25-20+5}

\small \boldsymbol{f(g(5))=6-\sqrt{10}}